For each of the following pairs of functions, determine whether 𝒇(𝒏) = 𝑢(π’ˆ(𝒏)) or


π’ˆ(𝒏) = 𝑢(𝒇(𝒏)).


a. 𝑓(𝑛) = 𝑛(𝑛 βˆ’ 1)⁄2 and 𝑔(𝑛) = 6𝑛


b. 𝑓(𝑛) = 𝑛 + 2βˆšπ‘› and 𝑔(𝑛) = 𝑛^2


c. 𝑓(𝑛) = 𝑛 + log 𝑛 and 𝑔(𝑛) = π‘›βˆšπ‘›


d. 𝑓(𝑛) = 𝑛 log 𝑛 and 𝑔(𝑛) = π‘›βˆšπ‘›/2


e. 𝑓(𝑛) = 2(log 𝑛)^2


and 𝑔(𝑛) = log 𝑛 + 1


1
Expert's answer
2021-12-21T12:26:28-0500

a.

"g(n)=6n\le 6f(n)=6n(n-1)"

π’ˆ(𝒏) = 𝑢(𝒇(𝒏))


b.

"𝑓(𝑛) = 𝑛 + 2\sqrt𝑛\le 3n\le 3𝑔(𝑛) = 3𝑛^2"

𝒇(𝒏) = 𝑢(π’ˆ(𝒏))


c.

"𝑓(𝑛) = 𝑛 + log 𝑛\le 2n\le 2𝑔(𝑛) = 2𝑛\sqrt 𝑛"

𝒇(𝒏) = 𝑢(π’ˆ(𝒏))


e.

"𝑔(𝑛) = log 𝑛 + 1\le 2logn\le 2f(n)=2(log 𝑛)^2"

π’ˆ(𝒏) = 𝑢(𝒇(𝒏))


d.

"𝑓(𝑛) = 𝑛 log 𝑛\le n\sqrt n=2𝑔(𝑛)"

𝒇(𝒏) = 𝑢(π’ˆ(𝒏))